OpenStudy (anonymous):

Help me please 3.2sin2x=2+cosx in the interval [0,2pi] 4. tan3x(tanx-1)=0 5.2sin^2x+3sinx+1 in the interval [0,2pi}

OpenStudy (anonymous):

OUCH! These take long haha ill take just one! Which one you want...the most? Lol

OpenStudy (anonymous):

i need all, but i will take the first one .

OpenStudy (anonymous):

Alrighty then you have: \[\eqalign{ &5\cos(x)=\sqrt{3}+3\cos(x);x\in[0,2\pi] \\ &2\cos(x)=\sqrt{3} \\ &\cos(x)=\frac{\sqrt{3}}{2} }\] Now according to the CAST rule (let me know if you haven't learned that yet), cosin is positive in the first and fourth quadrants. By the way, do you want this in radians or degrees?

OpenStudy (anonymous):

radians

myininaya (myininaya):

You can use the unit circle to find when cos is sqrt(3)/2 @JustTooMuchh

myininaya (myininaya):

Use that circle to find when the x-coordinate is sqrt(3)/2

OpenStudy (anonymous):

k thanks